# Residual lifetime prediction for heterogeneous degradation data by Bayesian semi-parametric method

### v1

@@ 2.1 Bayesian semi-parametric degradation model
$G\sim DP(\gamma,G_{0})$
$\alpha\sim\pi_{1}(\alpha)$
$\sigma_{\epsilon}^{2}\sim\pi_{2}(\sigma_{\epsilon}^{2})$
Note that the parameter $\gamma$ controls the concentration of the prior for G about $G_{0}$. It can be proved that, for any measurable subset $A$ of $\Theta$, $E[G(A)]={G_{0}}(A)$ and $Var[G(A)]=\frac{{G_{0}}(A)[1-{G_{0}}(A)]}{1+\gamma}$. Observe that $\forall A\in\mathscr{B}$, $G(A)$ is highly concentrated about $G_{0}(A)$ for large values of $\gamma$. On the other hand, as $\gamma$ tends to zero the expected shape of G is different from the one assumed by the baseline probability measure $G_{0}$. Since $G$ is a discrete probability measure with probability 1, so there will be repetition among the drawn values from this distribution, in other words if we draw $m$ values say $\theta_{1},.....,\theta_{m}$ from $G$, then there will be $k_{m}$ unique values of $\theta_{i}$’s where $k_{m}\leq m$. One can observe that clustering is induced among $\beta_{i}$’s, $i=1,...,(n+1)$ due to the fact that there are similar values among the unobserved random element $\theta_{i}$’s. So there is a positive probability that given $\theta_{i}$’s the distribution of random effect $\beta_{i}$’s are same for distinct $i$’s. The expected number of clusters or, equivalently, the number of distinct $\theta_{i}$ among the $m$ drawn sample values is given by $E[k_{m}]=\sum_{i=1}^{m}\frac{\gamma}{\gamma+i-1}$. If $m\rightarrow\infty$, it follows that $E[k_{m}]\approx\gamma\log(\frac{\gamma+m}{\gamma})$. Our next goal is to choose a suitable expression for $f(\cdot),\pi_{1}(\cdot)$ and $\pi_{2}(\cdot)$ and construct the hierarchy of the model.
@@ 2.2 Dirichlet process mixture of normal distribution for degradation model
As discussed in Section 2.2, we assume that the mixing probability measure $G$ has a Dirichlet process prior denoted by $G\sim DP(\gamma,G_{0})$. Sethuraman [32] showed that $G$ can be represented as, $G(\cdot)=\sum\limits_{h=1}^{\infty}p_{h}\delta_{m_{h}}(\cdot)$, where $m_{h}$ are random samples from $G_{0}$, $0\leq p_{h}\leq 1$ $\forall$ $h$ and $p_{1}=V_{1}$, $p_{h}=(1-V_{1})(1-V_{2})...(1-V_{h-1})V_{h}$ $\forall$ $h\geq 2$. Ishawarn and Zarepour [25] represented this sum of infinitely many terms by sum of finitely many $N$ terms. This method approximate Dirichlet process prior by $\mathbb{G}_{N}$, where $\mathbb{G}_{N}(\cdot)=\sum\limits_{h=1}^{N}p_{h}\delta_{Z_{h}}(\cdot)$ and $\delta_{Z}(\cdot)$ denotes a discrete measure concentrated at $Z$. Note that $\mathbb{G}_{N}$ is a random probability measure, and as $N\rightarrow\infty$, it converges almost surely to a Dirichlet process with baseline probability measure $G_{0}$ and Dirichlet mass parameter $\gamma$, denoted by $DP(\gamma,G_{0})$. Let us consider, $\textbf{Z}=(Z_{1},..,Z_{N})$, and $Z_{1},....,Z_{N}$ are i.i.d random variables with distribution $G_{0}$ and also they are independent of $\textbf{p}=(p_{1},...,p_{N})$. A probability measure simulated from the prior $\mathbb{G}_{N}$ is defined by choosing its random weights $p_{1},...,p_{N}$ by the stick-breaking construction, where $p_{1}=V_{1}$, $p_{h}=(1-V_{1})(1-V_{2})...(1-V_{h-1})V_{h},\hskip 2.27621pth=2,...,N$, with $V_{1},V_{2},...V_{N-1}$ are i.i.d $Beta(1,\gamma)$ random variables, setting $V_{N}=1$ ensures that $\sum\limits_{h=1}^{N}p_{h}=1$.
@@
Ishawaran and Zarepour [25] utilized approximated Dirichlet process $\mathbb{G}_{N}$ and introduced block Gibbs sampler to do direct inference for $\mathbb{G}_{N}$. They have constructed an efficient MCMC method, which recast the nonparametric hierarchical model completely in terms of random variables. Let $K_{1},......,K_{n+1}$ be the classification variables of the random effects $\beta_{1},......,\beta_{n+1}$ and $K_{i}$’s are conditionally independent random variables given p which identifies $Z_{k}$ with each associated $\theta_{i}$, particularly $\theta_{i}=Z_{K_{i}}$, $i=1,...,(n+1)$. The clustering nature of the hidden variables $\theta_{i}$’s are described by the classification variables $K_{1},......,K_{n+1}$. Let us consider $\pi_{3}(\cdot)$ and $\pi_{4}(\cdot)$ be the prior distributions for p and Z respectively. It follows that the hierarchy for distribution of $\beta_{i}$. $i=1,....,(n+1)$ as presented in equation (4) can be rewritten as (See Ishawaran and Zarepour [25])
$\beta_{i}|\textbf{Z},K_{i}\sim f(\beta_{i}|Z_{K_{i}})$ (5)
$K_{i}|\textbf{p}\sim\sum\limits_{h=1}^{N}p_{h}\delta_{h}(\cdot)$
$\sim\pi_{3}(\textbf{p})$ p )
$\sim\pi_{4}(\textbf{Z})$ Z )
This representation also helps to the fact that any degradation trajectory that is significantly different from the others can be classified as a distinct cluster. In this work we have taken the truncation parameter $N$ as the total number of units on which the proposed degradation model is built.
@@
We consider a normal distribution as a prior for the fixed effect parameter $\alpha$ and hierarchical priors are considered for its parameters. To complete the prior specification of $\mathbb{G}_{N}(\cdot)$, we consider $Z_{h}=(\mu_{h},\sigma_{h}^{2})$, where $\mu_{h}$ and $\sigma_{h}^{2}$ are mean and variance of conditional distribution of random effect given $\bm{\mu},\bm{\sigma}^{2},K_{i}=h$ which is normal and denoted as $\mathcal{N}(\cdot)$ in this work. Normal prior is considered for $\mu_{h}$’s and gamma prior for $(\sigma_{h}^{2})^{-1}$, $h=1,....,N$ which we denote as $\mathcal{G}(\cdot)$. We consider Gamma distribution with the form $\mathcal{G}(a_{0},b_{0})$ so that mean is $\frac{a_{0}}{b_{0}}$. Gamma prior is considered for inverse of variance of error measurements $\sigma_{\epsilon}^{2}$. We also assume that $\mu_{h}\gg\sigma_{h}$, $h=1,...,N$. We discuss in details on the prior specification in Section 4. Observe that, classification variables of random effects $K_{i}$’s, $i=1,...,(n+1)$ are conditionally independent random variables given p, where p is constructed by stick-breaking construction. Note that $K_{n+1}$ is the classification variable corresponding to random effect of the new unit. The hierarchy of the model is given below :
[…]
@@
$\hskip 113.81102pt\mu_{\alpha}\sim\mathcal{N}(\mu_{1},\sigma_{1}^{2})$
$\hskip 108.12047pt(\sigma_{\alpha}^{2})^{-1}\sim\mathcal{G}(a_{1},a_{2})$
$\hskip 99.58464pt\beta_{i}|\bm{\mu},\bm{\sigma}^{2},K_{i}\sim\mathcal{N}(\mu_{K_{i}},{\sigma}_{K_{i}}^{2}),\hskip 5.69054pti=1,......,n+1$
$\hskip 99.58464ptK_{i}|p\sim\sum\limits_{h=1}^{N}p_{h}\delta_{h}(\cdot),\hskip 5.69054pti=1,......,n+1$
$\hskip 71.13188ptp_{1}=V_{1},\hskip 5.69054ptp_{h}=V_{h}\prod\limits_{j=1}^{h-1}(1-V_{j}),\hskip 5.69054pth=2,...,N,\hskip 5.69054ptV_{l}|\gamma\overset{\mathrm{i.i.d}}{\sim}Beta(1,\gamma),$
$\hskip 71.13188pt\hskip 5.69054ptl=1,....,N-1\hskip 5.69054pt\&\hskip 5.69054ptV_{N}=1$
$\hskip 113.81102pt\mu_{h}|\sigma_{z}^{2}\sim\mathcal{N}(m_{\mu},\sigma_{z}^{2}),\hskip 5.69054pth=1,......,N$
@@
$\hskip 113.81102pt({\sigma_{h}^{2}})^{-1}\sim\mathcal{G}(a,b),\hskip 5.69054pth=1,......,N$
@@
The posterior distribution of the parameters is given by
$\pi(\bm{\beta},\alpha,\mu_{\alpha},\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma|\bm{Y}=\bm{y}),$ (7)
where $\bm{\mu}=(\mu_{1},.......,\mu_{N})$, $\bm{\sigma}^{2}=(\sigma_{1}^{2},.......,\sigma_{N}^{2})$, $\bm{\beta}=(\beta_{1},.....,\beta_{n+1})$, $\textbf{K}=(K_{1},......,K_{n+1})$ and $\textbf{p}=(p_{1},......,p_{N})$. In the next section we consider derivation of residual lifetime distribution of a unit conditioned on the observed degradation measurements. We show that this distribution depends on the samples from posterior distribution of parameters and hence it is required to generate samples from the above-mentioned posterior distribution presented in (7).
@@ 3 Residual lifetime distribution
Our goal is to derive distribution function of residual lifetime of a new unit for which we have degradation measurements at time points $t_{1},....,t_{k}$. Suppose the degradation observations of the new unit has not crossed threshold value $D$ till time $t_{k}$, that is $\underset{1\leq a\leq k}{max}\{y_{new,t_{a}}\}$ $\leq D$. The residual lifetime $T_{new}$ of new unit functioning at time $t_{k}$ is defined to be the first time when the degradation measurement reaches the threshold value $D$ subtracted from $t_{k}$, which is expressed as follows.
@@
Next we derive the distribution of $T_{new}$. We make the following assumptions for deriving the residual lifetime distribution of a functioning unit.
@@
The distribution function of $T_{new}$ is obtained as.
$F_{T_{new}}(t)=$ $P[T_{new}\leq t]$ (9)
$=$ $1-P[T_{new}>t]$
$=$ $1-P[\underset{s\in(0,\infty)}{\text{inf}}(s:Y_{new,s+t_{k}}\geq D)>t]$
$=$ $2\displaystyle 1-P[Y_{new,t+t_{k}}<D],\hskip 5.69054ptby\hskip 5.69054pt\text{assumption}\hskip 2.27621pt$ 2
$=$ $P[Y_{new,t+t_{k}}\geq D].$
We compute distribution function of random variable $T_{new}$ given $\bm{Y}=\bm{y}$, which is denoted as $F_{T_{new}|\bm{Y}=\bm{y}}(t)$ where $\bm{y}$ is the observed value of the random variable $\bm{Y}$. This implies that it is enough to derive $P[Y_{new,t+t_{k}}\geq D|\bm{Y}=\bm{y}]$, since the event $[T_{new}\leq t]$ equals the event $[Y_{new,t+t_{k}}\geq D]$ under the assumption 2. Thus it is also possible to derive residual lifetime distribution of each unit of the training set by considering the individual unit as the unit for which residual lifetime is to be predicted and rest of the units as training set. So, if there are $n$ units in training set, then we will fix one unit as new unit and use $(n-1)$ units as training data. This process can be done for all $n$ units. In this work we consider that failure of a unit occurs when its observed degradation measurement crosses a predetermined threshold value say $D$. Sometimes degradation measurements of a unit are available after its degradation path crosses the threshold value. In our proposed method these information are also used to compute residual lifetime distribution. The residual lifetime distribution of a new unit conditioned on $\bm{Y}=\bm{y}$, is given by:
$F_{T_{new}|\bm{Y}=\bm{y}}(t)=$ $P[T_{new}\leq t|\bm{Y}=\bm{y}]$
@@
$=$ $\int_{\bm{\gamma_{1}}}P[Y_{new,t+t_{k}}\geq D,\bm{\gamma_{1}}|\bm{Y}=\bm{y}]d\bm{\gamma_{1}},\hskip 2.84526pt\bm{\gamma_{1}}=(\alpha,\beta_{1},....,\beta_{n},\beta_{new},\sigma_{\epsilon}^{2})$
@@
[…]
$=$ $E_{\bm{\gamma_{1}}|\bm{Y}=\bm{y}}\big{[}P[Y_{new,t+t_{k}}>D|\bm{\gamma_{1}}]\big{]}.\hskip 176.40707pt(10)$
It may be noted that $F_{T_{new}|\bm{Y}=\bm{y}}(t)$ can not be obtained analytically. We compute it approximately based on observations generated from the posterior distribution of $\bm{\gamma_{1}}$. Let $\alpha^{(i)}$, $\beta_{1}^{(i)}$,…., $\beta_{n}^{(i)}$, $\beta_{new}^{(i)}$ and $(\sigma_{\epsilon}^{2})^{(i)}$, $i=1,.....,N_{1}$ are the samples generated from posterior distribution of $\bm{\gamma_{1}}$ and $N_{1}$ is number of posterior samples. We denote this approximated distribution function as $F_{T_{new}|\bm{Y}=\bm{y}}^{approx}(t)$, which is given as follows.
$F_{T_{new}|\bm{Y}=\bm{y}}^{approx}(t)$ $=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[Y_{new,t+t_{k}}\geq D|\bm{\gamma_{1}}=\bm{\gamma_{1}}^{(i)}]$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[Y_{new,t+t_{k}}\geq D|\alpha=\alpha^{(i)},\beta_{1}=\beta_{1}^{(i)},...,\beta_{n}=\beta_{n}^{(i)},\beta_{new}=\beta_{new}^{(i)},\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]$

### v2

@@ 2.1 Bayesian semi-parametric degradation model
$G\sim DP(\gamma,G_{0}),\alpha\sim\pi_{1}(\alpha),\sigma_{\epsilon}^{2}\sim\pi_{2}(\sigma_{\epsilon}^{2})$
The hierarchical representation given by model (4), explains the semi-parametric nature of the proposed model. We also assume that the random variable for degradation measurement at $t$th time, $Y_{it}$ are conditionally independent with $\beta_{1},\dots,\beta_{i-1},\beta_{i+1},...,\beta_{n+1}$ given $\alpha,\beta_{i}$ and $\sigma_{\epsilon}^{2}$, $i=1,\dots,(n+1)$, and $\forall\hskip 2.84526ptt>0$. The parameter $\gamma$ controls the concentration of the prior for G about $G_{0}$. It can be proved that, for any measurable subset $A$ of $\Theta$, $E[G(A)]={G_{0}}(A)$ and $Var[G(A)]=\frac{{G_{0}}(A)[1-{G_{0}}(A)]}{1+\gamma}$. Observe that $\forall A\in\mathscr{B}$, $G(A)$ is highly concentrated about $G_{0}(A)$ for large values of $\gamma$. On the other hand, as $\gamma$ tends to zero the expected shape of G is different from the one assumed by the baseline probability measure $G_{0}$. Since $G$ is a discrete probability measure with probability 1, so there will be repetition among the drawn values from this distribution, in other words if we draw $m$ values say $\theta_{1},\dots,\theta_{m}$ from $G$, then there will be $k_{m}$ unique values of $\theta_{i}$’s where $k_{m}\leq m$. One can observe that clustering is induced among $\beta_{i}$’s, $i=1,\dots,(n+1)$ due to the fact that there are similar values among the unobserved random element $\theta_{i}$’s. So there is a positive probability that given $\theta_{i}$’s the distribution of random effect $\beta_{i}$’s are same for distinct $i$’s. The expected number of clusters or, equivalently, expected number of distinct $\theta_{i}$ among the $m$ drawn sample values is given by $E[k_{m}]=\sum_{i=1}^{m}\frac{\gamma}{\gamma+i-1}$. If $m\rightarrow\infty$, it follows that $E[k_{m}]\approx\gamma\log(\frac{\gamma+m}{\gamma})$. Our next goal is to choose a suitable expression for $f(\cdot),\pi_{1}(\cdot)$ and $\pi_{2}(\cdot)$ and construct the hierarchy of the model.
@@ 2.2 Dirichlet process mixture of normal distribution for degradation model
As discussed in Section 2.1, we assume that the mixing probability measure $G$ has a Dirichlet process prior denoted by $G\sim DP(\gamma,G_{0})$. Sethuraman [50] showed that $G$ can be represented as, $G(\cdot)=\sum\limits_{h=1}^{\infty}p_{h}\delta_{m_{h}}(\cdot)$, where $m_{h}$ are random samples from $G_{0}$, $0\leq p_{h}\leq 1$ $\forall$ $h$ and $p_{1}=V_{1}$, $p_{h}=(1-V_{1})(1-V_{2})\dots(1-V_{h-1})V_{h}$ $\forall$ $h\geq 2$. Ishawarn and Zarepour [37] represented this sum of infinitely many terms by sum of finitely many $N$ terms. This method approximate Dirichlet process prior by $\mathbb{G}_{N}$, where $\mathbb{G}_{N}(\cdot)=\sum\limits_{h=1}^{N}p_{h}\delta_{Z_{h}}(\cdot)$ and $\delta_{Z}(\cdot)$ denotes a discrete measure concentrated at $Z$. Note that $\mathbb{G}_{N}$ is a random probability measure, and as $N\rightarrow\infty$, it converges almost surely to a Dirichlet process with baseline probability measure $G_{0}$ and Dirichlet mass parameter $\gamma$, denoted by $DP(\gamma,G_{0})$. Let us consider, $\textbf{Z}=(Z_{1},\dots,Z_{N})$, and $Z_{1},\dots,Z_{N}$ are i.i.d random variables with absolutely continuous distribution function $G_{0}$ and also they are independent of $\textbf{p}=(p_{1},\dots,p_{N})$. A probability measure simulated from the prior $\mathbb{G}_{N}$ is defined by choosing its random weights $p_{1},\dots,p_{N}$ by the stick-breaking construction, where $p_{1}=V_{1}$, $p_{h}=(1-V_{1})(1-V_{2})\dots(1-V_{h-1})V_{h},\hskip 2.27621pth=2,\dots,N$, with $V_{1},V_{2},\dots,V_{N-1}$ are i.i.d $Beta(1,\gamma)$ random variables, setting $V_{N}=1$ ensures that $\sum\limits_{h=1}^{N}p_{h}=1$.
@@
Ishawaran and Zarepour [37] utilized this approximated Dirichlet process $\mathbb{G}_{N}$ and introduced block Gibbs sampler to estimate posterior $\mathbb{G}_{N}^{*}$. They have constructed an efficient MCMC method, which represents the nonparametric hierarchical model completely in terms of random variables. We use this technique to represent the hierarchical structure of the random effect distribution, given by model (4). Let us consider, $K_{1},\dots,K_{n+1}$ be the classification variables of the random effects $\beta_{1},\dots,\beta_{n+1}$ and $K_{i}$’s are conditionally independent random variables given p which identifies $Z_{k}$ with each associated $\theta_{i}$, particularly $\theta_{i}=Z_{K_{i}}$, $i=1,\dots,(n+1)$. The clustering nature of the hidden variables $\theta_{i}$’s are described by the classification variables $K_{1},\dots,K_{n+1}$. Let us consider $\pi_{3}(\cdot)$ and $\pi_{4}(\cdot)$ be the prior distributions for p and Z respectively. It follows that the hierarchy for distribution of $\beta_{i}$. $i=1,\dots,(n+1)$ as presented in equation (4) can be rewritten as (See Ishawaran and Zarepour [37])
$\beta_{i}|\textbf{Z},K_{i}\overset{\mathrm{ind}}{\sim}f(\beta_{i}|Z_{K_{i}})$ (5)
$K_{i}|\textbf{p}\overset{\mathrm{iid}}{\sim}\sum\limits_{h=1}^{N}p_{h}\delta_{h}(\cdot)$
$\textbf{p}\sim\pi_{3}(\textbf{p}),\textbf{Z}\sim\pi_{4}(\textbf{Z})$
@@
The expression given by model (5) implies that a random effects on distinct units can have distribution with identical parameter values which is induced due to discrete nature of classification random variables. This representation helps to the fact that any degradation trajectory that is significantly different from the others can be classified as a distinct cluster. In this work we take the truncation parameter $N$ as the total number of units on which the proposed degradation model is built. We consider a normal distribution as a prior for the fixed effect parameter $\alpha$ and hierarchical priors are considered for its parameters. To complete the prior specification of $\mathbb{G}_{N}(\cdot)$, we consider $Z_{h}=(\mu_{h},\sigma_{h}^{2})$, where $\mu_{h}$ and $\sigma_{h}^{2}$ are mean and variance of conditional distribution of random effect given $\bm{\mu},\bm{\sigma}^{2},K_{i}=h$ which is normal and denoted as $\mathcal{N}(\cdot)$ in this work. Normal prior is considered for $\mu_{h}$’s and gamma prior for $(\sigma_{h}^{2})^{-1}$, $h=1,\dots,N$ which we denote as $\mathcal{G}(\cdot)$. We consider Gamma distribution with the form $\mathcal{G}(a_{0},b_{0})$ so that mean is $\frac{a_{0}}{b_{0}}$. Gamma prior is considered for inverse of variance of error measurements $\sigma_{\epsilon}^{2}$. The hierarchical structure of the proposed model is as follows,
$\hskip 71.13188ptY_{ij}|\alpha,\beta_{i},\sigma_{\epsilon}^{2}\overset{\mathrm{ind}}{\sim}\mathcal{N}(\alpha+\beta_{i}t_{ij},\sigma_{\epsilon}^{2}),\hskip 5.69054pti=1,\dots,n+1,j=1,\dots,n_{i}$ (6)
@@
$\hskip 99.58464pt\beta_{i}|\bm{\mu},\bm{\sigma}^{2},K_{i}\overset{\mathrm{ind}}{\sim}\mathcal{N}(\mu_{K_{i}},{\sigma}_{K_{i}}^{2}),\hskip 5.69054pti=1,\dots,n+1$
$\hskip 99.58464ptK_{i}|p\overset{\mathrm{iid}}{\sim}\sum\limits_{h=1}^{N}p_{h}\delta_{h}(\cdot),\hskip 5.69054pti=1,\dots,n+1$
$\hskip 71.13188ptp_{1}=V_{1},\hskip 5.69054ptp_{h}=V_{h}\prod\limits_{j=1}^{h-1}(1-V_{j}),\hskip 5.69054pth=2,\dots,N,\hskip 5.69054ptV_{l}|\gamma\overset{\mathrm{i.i.d}}{\sim}Beta(1,\gamma),$
$\hskip 71.13188pt\hskip 5.69054ptl=1,\dots,N-1\hskip 5.69054pt\&\hskip 5.69054ptV_{N}=1$
$\hskip 113.81102pt\mu_{h}|\sigma_{z}^{2}\sim\mathcal{N}(m_{\mu},\sigma_{z}^{2}),\hskip 5.69054pth=1,\dots,N$
@@
$\hskip 113.81102pt({\sigma_{h}^{2}})^{-1}\sim\mathcal{G}(a,b),\hskip 5.69054pth=1,\dots,N$
@@
In this work, we use model (6) to derive the posterior distribution of parameters. Observe that, classification variables of random effects $K_{i}$’s, $i=1,\dots,(n+1)$ are conditionally independent random variables given p, where p is constructed by stick-breaking construction. Note that $K_{n+1}$ is the classification variable corresponding to random effect of the new unit. The posterior distribution of the parameters is given by
$\pi(\bm{\beta},\alpha,\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma|\bm{Y}=\bm{y}),$ (7)
where $\bm{\mu}=(\mu_{1},\dots,\mu_{N})$, $\bm{\sigma}^{2}=(\sigma_{1}^{2},\dots,\sigma_{N}^{2})$, $\bm{\beta}=(\beta_{1},\dots,\beta_{n+1})$, $\textbf{K}=(K_{1},\dots,K_{n+1})$ and $\textbf{p}=(p_{1},\dots,p_{N})$. In the next section we consider derivation of residual lifetime distribution of a unit conditioned on the observed degradation measurements. We show that this distribution depends on the samples from posterior distribution of parameters and hence it is required to generate samples from the posterior distribution given by expression (7). We discuss in details on the prior specification in Section 4.
@@ 3 Residual lifetime distribution
Our goal is to derive distribution function of residual lifetime of a new unit for which we have degradation measurements at time points $t_{1},\dots,t_{k}$. In this work, we consider that failure of a unit occurs when the random variable for observed degradation measurement crosses a predetermined threshold value $D$. Suppose the degradation observations of the new unit has not crossed threshold value $D$ till time $t_{k}$, that is $\underset{1\leq a\leq k}{max}\{y_{new,t_{a}}\}$ $\leq D$. Let $T_{new}$ be a random variable which is defined as the first time the degradation measurement reaches the threshold value $D$ subtracted from $t_{k}$, which is given as follows in equation (8).
@@
The residual lifetime of the new unit functioning at time $t_{k}$ is defined as the first time the degradation measures crosses the threshold value $D$ given that the degradation measures at time point $t_{1},\dots,t_{k}$ are less than $D$. So, given the collection of random variables for degradation observations $\textbf{Y}_{new}$, it is enough to to derive the conditional distribution function of random variable $T_{new}$ given $\textbf{Y}_{new}=\textbf{y}_{new}$. We make the following assumptions for deriving the residual lifetime distribution of a functioning unit.
@@
The distribution function of $T_{new}$ given $\textbf{Y}_{new}=\textbf{y}_{new}$ is obtained as.
$F_{T_{new}|\textbf{Y}_{new}}(t)=$ $P[T_{new}\leq t|\textbf{Y}_{new}=\textbf{y}_{new}]$ (9)
$=$ $1-P[T_{new}>t|\textbf{Y}_{new}=\textbf{y}_{new}]$
$=$ $1-P[\underset{s\in(0,\infty)}{\text{inf}}(s:Y_{new,s+t_{k}}\geq D)>t|\textbf{Y}_{new}=\textbf{y}_{new}]$
$=$ $1-P[Y_{new,t+t_{k}}<D|\textbf{Y}_{new}=\textbf{y}_{new}],\hskip 5.69054pt\text{by}\hskip 5.69054pt\text{assumption}\hskip 2.27621pt2$
$=$ $P[Y_{new,t+t_{k}}\geq D|\textbf{Y}_{new}=\textbf{y}_{new}].$
The degradation model as defined by model (6), implies that the distribution function of residual lifetime of the new unit depends on the parameters $\alpha,\beta_{new}$ and $\sigma_{\epsilon}^{2}$. So, given these parameters and individual degradation observations, the distribution function of $T_{new}$ is as follows.
$F_{T_{new}|\textbf{Y}_{new},\alpha,\beta_{new},\sigma_{\epsilon}^{2}}(t)=$ $P[T_{new}\leq t|\textbf{Y}_{new}=\textbf{y}_{new},\alpha=\alpha^{*},\beta_{new}=\beta_{new}^{*},\sigma_{\epsilon}^{2}=\sigma_{\epsilon}^{2*}]$
$=$ $1-P[T_{new}>t|\textbf{Y}_{new}=\textbf{y}_{new},\alpha=\alpha^{*},\beta_{new}=\beta_{new}^{*},\sigma_{\epsilon}^{2}=\sigma_{\epsilon}^{2*}]$
$=$ $P[Y_{new,t+t_{k}}\geq D|\alpha=\alpha^{*},\beta_{new}=\beta_{new}^{*},\sigma_{\epsilon}^{2}=\sigma_{\epsilon}^{2*}].$
$\hskip 5.69054pt[\text{using (6)}]$
[…]
Observe that the expression in equation (10) is a distribution function when $\beta_{new}^{*}>0$. If we use equation (9) for computing the distribution function of residual lifetime of the new unit, only individual degradation observations of this new unit are utilized, whereas it would be more appropriate if one can use the degradation observations of $(n+1)$ units. In order to tackle this issue, we compute distribution function of the random variable $T_{new}$ given $\bm{Y}=\bm{y}$, which is denoted as $F_{T_{new}|\bm{Y}}(t)$ where $\bm{y}$ is the observed value of the random vector $\bm{Y}$ and $\bm{Y}=[\textbf{Y}_{train},\textbf{Y}_{new}]$. So it is enough to derive $P[Y_{new,t+t_{k}}\geq D|\bm{Y}=\bm{y}]$, since the event $T_{new}\leq t]$ equals the event $Y_{new,t+t_{k}}\geq D]$ under the assumption 2. Thus, it is also possible to derive residual lifetime distribution of each unit of the training set by considering the individual unit as the unit for which residual lifetime is to be predicted and rest of the units as training set. So, if there are $n$ units in training set, then we will fix one unit as a new unit and use (n1)(n-1) units as training data. This process can be done for all $n$ units.
Sometimes degradation measurements of a unit are available after its degradation path crosses the threshold value. In our proposed method these information are also used to compute residual lifetime distribution. The residual lifetime distribution of a new unit conditioned on $\bm{Y}=\bm{y}$, is given by,
$F_{T_{new}|\bm{Y}}(t)=$ $P[T_{new}\leq t|\bm{Y}=\bm{y}]$
@@
$=$ $\int_{\bm{\gamma_{1}}}P[Y_{new,t+t_{k}}\geq D,\bm{\gamma_{1}}|\bm{Y}=\bm{y}]d\bm{\gamma_{1}},\hskip 2.84526pt\bm{\gamma_{1}}=(\alpha,\beta_{1},\dots,\beta_{n},\beta_{new},\sigma_{\epsilon}^{2})$
$=$ $\int_{\bm{\gamma_{1}}}P[Y_{new,t+t_{k}}\geq D,|\bm{\gamma_{1}},\bm{Y}=\bm{y}]\pi(\bm{\gamma_{1}}|\bm{Y}=\bm{y})d\bm{\gamma_{1}},\hskip 2.84526pt,\hskip 2.84526ptd\bm{\gamma_{1}}=d\alpha d\beta_{1}\dots d\beta_{new}d\sigma_{\epsilon}^{2}$
$=$ $\int_{\bm{\gamma_{1}}}P[Y_{new,t+t_{k}}\geq D|\bm{\gamma_{1}}]\pi(\bm{\gamma_{1}}|\bm{Y}=\bm{y})d\bm{\gamma_{1}}$
@@
It may be noted that to obtain analytic expression of $F_{T_{new}|\bm{Y}}(t)$, represented in equation (11), a large number of multivariate integrals need to be computed. To avoid this issue, we compute this expression given in equation (11) approximately based on observations generated from the posterior distribution of $\bm{\gamma_{1}}$. Let $\alpha^{(i)}$, $\beta_{1}^{(i)}$, …, $\beta_{n}^{(i)}$, $\beta_{new}^{(i)}$ and $(\sigma_{\epsilon}^{2})^{(i)}$, $i=1,\dots,N_{1}$ are the samples generated from posterior distribution of $\bm{\gamma_{1}}$ and $N_{1}$ is number of posterior samples. Let us consider $\epsilon_{new}$ be the random variable for measurement error for the degradation measurement at the time point $(t+t_{k})$. Next, we derive the the approximated distribution function denoted as $F_{T_{new}|\bm{Y}}^{approx}(t)$, which is given as follows.
$F_{T_{new}|\bm{Y}}^{approx}(t)$ $=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[Y_{new,t+t_{k}}\geq D|\bm{\gamma_{1}}=\bm{\gamma_{1}}^{(i)}]$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[Y_{new,t+t_{k}}\geq D|\alpha=\alpha^{(i)},\beta_{1}=\beta_{1}^{(i)},\dots,\beta_{n}=\beta_{n}^{(i)},\beta_{new}=\beta_{new}^{(i)},\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]$